Tohban Report 2013-05-01

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 24 Apr 2013
End Date: 01 May 2013
Tohban: Mitsuo Oka
Tohban email:
Next Tohban: Lindsay Glesener
List all reports


Solar Activity

Solar activity was moderate this week. Many C class flares were produced by AR 11726 in the earlier half of the week and by AR 11730/11731 in the later half of the week. There are currently 6 active regions on the disk. AR 11731 is expected to remain active for the next few days and STEREO B shows one active region which would rotate onto the disk over the next few days.

How many GOES flares occurred?

 Flares above B, C, M, X class were      3    72     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      3    55     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      1    30     0     0

There were RHESSI flares/GOES flares 507 / 75 over the time range 23-Apr-13 30-Apr-13

Memory Management

The SSR was mostly emptying at the end of the daily passes with a maximum fill level of ~25%.

Spacecraft Status

The cold plate 1 and 2 are at 112 K and 110 K respectively. Spint rate is 14.6 rpm. No significant change in the cryocooler status.

Data Gaps

There was no data gap this week.

Detector issues

Other notes


On April 25, during passes at 03:30 UT and 05:30 UT, a rebalancing of the cryocooler was performed as a measure to lower the power consumption and the overall increase trend in temperature. However, increasing/decreasing of both amplitude and phase of the balancer only led to vibration increase. The amplitude and phase of the balancer were put back to original values.

Brian reported further changes in state-of-health plots. Since last week, the Sun fraction has increased to peak at 0.71 and the Cold Plate 2 temperature has decreased to 111 K. This is to be compared with 106.5 K at the time of the previous similar peak in the Sun fraction that occurred about Feb. 10. So there has been a 4.5 K increase in about 2.5 months or about 20 K per year. The accelerometer reading has stayed constant just below 16 mg for the last week or so.

Next course of action would be to rebalance with finer steps and to adjust both amplitude and phase simultaneously.

<Gamma Ray Burst>

RHESSI detected a large GRB event at 07:47 UT of April 27. An initial report can be found in the Nugget article by Andre Csillaghy and David Smith ([1]).

Spacecraft Management

Decimation Normal/vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/- 4 minutes
Night time data (rears) Taking data at nights
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? No
Personal tools